Solve for x.
3x^2= -17x-10

Answer by dabanfield(803)

3x^2= -17x-10
3x^2 + 17x + 10 = 0
(3x + 2)*(x + 5) = 0
Solve:
3x + 2 = 0
x + 5 = 0